OpenStudy (anonymous):

Under three years free Three to twelve years $36 Over twelve years $48 On a certain day, the company took 92 people on whale watching trips. There were 40 children aged twelve and under, of which some children were under three years. If x represents the number of children undr three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

Well, first let's see if we can make C mean something a tad more... useful. C=total money = $36*(#of children 3-12) + $48 (over 12). Since there were 92 people and 40 children, that means that 92-40= 52 of them were adults. This means that the equation above is now C=36*(#of children 3-12) +48*(52 adults)=36*(#of children 3-12) + 2496 Now, keep the above equation in mind. #of children 3-12 = 40 children - x children under 3 years old. You can plug this equation into the earlier equation so that: C= 36*(40 children - x children under 3) +2496. Solve for x: C = 1440 - 36x + 2496 36x = 1440 + 2496 - C = 3936 - C \[x=\frac{ 3936-C }{ 36 }\] Sorry if this is a bit confusing, I tend to make my explanations a tad complicated some times. And of course, you may want to check my math. I'm pretty confident in it, but I've learned never to say that I'm 100% confident ><